the sum of two numbers is 29 the difference between the same two numbers is 17.use a table to find the two numbers
1
Expert's answer
2012-03-20T08:09:56-0400
Let's denote two numbers as x,y
So we have
x+y=29&
x-y=17
Let's add two equations:
2x=29+17=46 whence x=23
y=29-x=29-23=6

the answer:& 23 and 6